# DualXDA: Towards Sparse, Efficient and Explainable Data Attribution in Large AI Models

### v2

@@ 6 Discussion & Conclusion
In this work, we introduce DualXDA, a framework that allows to scale interpretable data attribution to large AI models. DualXDA consists of two parts: First, DualDA is an efficient surrogate-based DA method for explaining AI models using the penultimate hidden representations to define a kernel map to fit a multiclass SVM. This strategy naturally provides global as well as local attributions as an approximate decomposition of the model output. Additionally, DualDA produces sparse attributions by virtue of the associated optimization process, which is beneficial for reducing the explanations’ complexity and volume. Second, XDA combines data attribution with feature attribution by mapping data attribution values back to the train and test input domain simultaneously, to identify relevant corresponding features and explain previously opaque data attribution results. Together, our framework overcomes the scaling problems of popular DA approaches and allows the user to gain additional insight due to its sparsity and feature-level explanations. Our contributions establish a foundation for explainable AI at unprecedented scale, enabling transparent and efficient analysis of large-scale neural architectures. DualXDA enables the development of interpretable and accountable AI systems by providing novel analytical capabilities that can be applied to state-of-the-art models in real-world applications beyond research settings due to its computational efficiency.
@@
Finally, we employ XDA to gain deeper insight into the DualDA attributions across three different datasets. Through qualitative analysis, we demonstrate how XDA explains nuances in model predictions and confounding factors in training data, while showing robustness in input ablation studies. To our knowledge, this is the first work to propose a methodology to explain the interaction of training and test datapoints from both perspectives, on feature level. By unifying the attributions in two orthogonal approaches in XAI, feature attribution and data attribution, we obtain a method that leverages the benefits of both approaches, constituting a novel explanation paradigm.
DualXDA, despite achieving substantially greater computational efficiency than competing approaches, still presents certain limitations primarily associated with the SVM training procedure. Standard SVM training lacks parallelization capabilities and exhibits computational complexity that scales with the size of the dataset, potentially creating bottlenecks for applications with large training sets. Nevertheless, in our experience early stopping before convergence produces effective results with reduced computational cost. Further acceleration can be achieved using GPU-based implementations such as thunderSVM (Wen et al., 2018) or modern parallelizable SVM training algorithms (Chang et al., 2012).
We see three potential avenues for extending this research as part of future work: Firstly, the DualXDA framework can be extended to different variants of the classical Support Vector Machine: Support Vector Regression (Drucker et al., 1996) may be used as a surrogate for the final layer of deep regression models. With the use of One-Class SVMs (Schölkopf et al., 2001), we can understand why a test point is considered an outlier using the training data. Incremental SVMs (Cauwenberghs & Poggio, 2000) may provide an efficient tool for implementing DualXDA in active learning settings. Finally, Relevance Vector Machines (Tipping, 2001) use a Bayesian approach by deploying an Automatic Relevance Determination prior, which allows for a probabilistic classification and may help to select support vectors which are further away from the margin and thus more representative of the data – it remains to be seen whether this is a desirable property for the tasks associated with DA. Secondly, prior research suggests that combining DA methods over an ensemble of models significantly improves its efficacy (Park et al., 2023). If this trend holds true, it provides a simple way to further boost the performance of our method. Finally, as DualXDA is designed for increased computational efficiency in order to be applicable for current state-of-the-art models with a large parameter space such as Large Language Models, we are interested in applications of DualXDA to these models to better understand how they are influenced by their training data. The value of this line of research lies in the potential to provide a method for the energy efficient interpretation of large-scale models with implications towards many real world applications, such as training dataset distillation for more efficient model training.
@@ Appendix A Derivation of DualDA using the SVM framework
Given the soft-margin optimization problem of (Crammer & Singer, 2001)
@@
To enforce the stationarity condition (Kuhn & Tucker, 1951), we take derivatives with respect to the primal variables and set them to 0:
@@
To solve this problem, we first derive the dual problem by plugging Equation A.8 in the Lagrangian and maximize it subject to the dual feasibility and complementary slackness conditions (Kuhn & Tucker, 1951):
We refer the reader to (Crammer & Singer, 2001) for the derivations.
@@ Appendix C Derivation of DualDA using the Influence Functions Framework – Proof of Equation 8
For a given test point with features and an arbitrary class , our objective is to determine how the class score responds to the infinitesimal downweighting of the -th datapoint. Using Corollary 1 in (Naujoks et al., 2024), we derive
@@ D.1 Data Attribution Methods
As DualDA is already defined in Section 3, we now specify the other DA methods we compare our approoach to in this paper. In all experiments, we employ GPU implementations and follow the best practices as suggested by original authors of the papers or code bases, as explained in the individual sections. This includes the use of random projections and CUDA-level implementations to efficiently perform these projections for TRAK. For the caching phase of DualDA, the extraction of latent features is run using GPU capabilities, while the training of the SVM uses CPU. The sequential nature of the underlying SVM optimization algorithm does not allow for a straightforward GPU parallelization of existing implementations (Keerthi et al., 2008; Chiang et al., 2016).
@@ D.1.1 Influence Functions
Influence Functions (IF) originates from the field of robust statistics, where it was initially used for the analysis of the effects of outliers on linear models (Srikantan, 1961; Hammoudeh & Lowd, 2022). Koh and Liang (Koh & Liang, 2017) proposed using this method for local data attribution of deep neural networks. It approximates the effect of discarding a training point using a second-order Taylor approximation. However, this approximation is conditioned on the current trained parameters of the network and hence includes information about the decision making strategies employed by the trained model. Data attribution by IF is given by
@@
Computation of the inverse Hessian is computationally prohibitive for modern architectures with millions of parameters. For this reason, (Koh & Liang, 2017) propose to use the LiSSA algorithm (Agarwal et al., 2017) which approximates the inverse Hessian vector product for each test point using an iterative procedure. While this mitigates the computational cost, it does not entirely solve the problem: computing explanations for a single test point can still take a long time, possibly in the order of hours (Hammoudeh & Lowd, 2022). In our experiments, we have followed the authors’ suggestions to determine hyperparameters which produce meaningful influence values while minimizing the required computation time. First, we have used only the final layer of deep neural networks to compute gradients for influence estimation with LiSSA. Secondly, we have set the total number of iterations dedicated to the estimation of the inverse Hessian vector product to the training dataset size in each experiment.
@@
Finally, the authors of Grosse et al. (Grosse et al., 2023) propose to solve the inverse-Hessian-vector product using an Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC), which is derived from an approximation of the Fisher information of the network. As hyperparameters, we have used the default values given in the original code release.
@@ D.1.2 TracIn
In the implementation of these methods, we have employed random projections on gradients to be able to cache training gradients, following (Pruthi et al., 2020). This allowed the attributions to be achieved in a practical time scale. We have used 128 random projections.
@@ D.2 Model Training
For CIFAR-10, we train a ResNet-18 (He et al., 2016) with random cropping and flipping as data augmentations, with a learning rate of 0.0003. We further include a weight decay term with a coefficient of 0.01 as part of the loss term.
@@ E.1 Sanity Checks
Hanawa et al. (Hanawa et al., 2021) suggest two sanity checks to determine the quality of DA methods.
@@
Hanawa et al. further suggest that attributions should be able to detect different subpopulations in the same class (Hanawa et al., 2021). To create this condition, they propose using an arbitrary partitioning of classes, and group each set in the partition into a super group. In practice, we have selected pairs of classes to group. These pairs of classes are chosen to be dissimilar to ensure that different strategies are learned for samples belonging to different subclasses. After training a neural network on the modified dataset, a successful DA methods should attribute highly to training samples that belong to the same subclass as the original class of the test sample.
To express this condition formally, following (Hanawa et al., 2021), let be an oracle function that returns the true (sub)class of the datapoint given as its input. Similar to the Identical Class Test, we report
@@ E.2 Metrics Emulating Downstream Tasks
The following metric mimics the downstream task of identifying mislabeled samples in the training data (Koh & Liang, 2017; Yeh et al., 2018; Pruthi et al., 2020). Correctly labelled datapoints can be classified by a model by relying on similar training samples of the same class. In contrast, mislabeled samples diverge from the typical feature patterns of their labeled class and consequently require individual memorization rather than generalization. Therefore, mislabelled datapoints are highly influential on their own model prediction, allowing the user to leverage DA approaches to detect mislabeled datapoints. To simulate this scenario, we randomly change the labels of a subset of the training data and retrain the model. We then use the attribution method to calculate each training sample’s self-influence, . To quantify the performance of the mislabeling detection by a DA method, we traverse the training data set in descending order of self-influence and record the cumulative density function of the percentage of mislabeled samples we have identified. We report the area under the curve of this cumulative density function, linearly transformed to a score between 0 and 1, where 0 corresponds to the minimum achievable area, 1 represents the maximum achievable area, and random attribution by sampling from a uniform distribution yields an expected score of 0.5.
@@
Another candidate downstream application for DA is to identify the causes of shortcut learning and backdoor poisoning attacks. If we modify the training dataset to introduce an artificial shortcut to trigger a certain model prediction, DA methods should attribute the responsible training images highly whenever the shortcut feature is present in the test sample. Similarly to previous work (Koh & Liang, 2017; Hammoudeh & Lowd, 2022), we formulate a shortcut detection metric: determine a perturbation which takes a sample and adds a shortcut perturbation on it. Importantly, applies a consistent and detectable perturbation to all chosen samples. We use a small box in the center, and a frame around the images in our experiments.
We select a class as the shortcut class, and randomly sample a subset of datapoints from that class to apply . By retraining the model on this perturbed training data, the model learns a connection between the artifact and the class . We now create a particular dataset for evaluation by applying the perturbation to all samples and collecting all perturbed samples, which are not originally from class but who, after the perturbation is applied, are classified as class , i.e. . This ensures that if a test point is now classified as class , it is primarily due to the influence of the perturbation and therefore, the perturbed training samples should be most influential for the classification as class . For all such test points, we report the Area Under the Precision-Recall Curve (AUPRC) score for the detection of perturbed training samples, following (Hammoudeh & Lowd, 2022).
@@ E.3 Retraining-based Metrics
The counterfactual relevance of training data can be empirically assessed through model retraining under an identical training regime with modified training sets. As explained in Section 2.2, this approach is the motivation for methods based on Influence Functions (Koh & Liang, 2017). The subsequent metrics evaluate the correspondence between observed changes in model behaviour and the attributions of the training data.
@@
LDS measures the effectiveness of attributions as an estimator for the effect of leaving out training samples by retraining the model on general subsets of the training data (Park et al., 2023). Given attributions , and a subset of training data indices , we define
@@
A downstream application of data attribution is dataset distillation: limiting the training dataset to fewer examples, while maintaining high test accuracy (Guo et al., 2022; Joaquin et al., 2024). In contrast to LDS, we are not interested in the model performance for randomly chosen subsets of the training data, but instead for the highest attributed subset. To measure this in a metric, we need a strategy to determine the global importance of a training point for the model performance on the entire test set. To achieve this, we compute the average value of the absolute attributions of a training point over the test set. We then restrict the training data to the most globally relevant datapoints and measure the test loss of a model retrained on the restricted training set. We report the test loss when retrained on 10% of the data as well as a weighted average
@@
Inspired by previous work, which investigates the sensitivity of model predictions to ablation of the training data (Ilyas et al., 2022), we measure the test loss of a model that was retrained on a train set which had its most influential datapoints removed. Similarly to Coreset Selection, we consider the highest attributed subset. But in contrast, we aim to judge how much exclusive relevant information this subset holds that would induce an increase in the loss when the subset is removed. We use the same strategy as in Coreset Selection to determine the global importance of training samples. We report the test loss when retrained without the 10% most influential data as well as a weighted average

### v3

@@ 5.2 Ablation Study
6 Limitations & Future Work
The current formulation of DualXDA employs a multiclass SVM as a surrogate to achieve Data Attribution, which restricts the demonstrated applications to classification settings. However, the framework does not rely on this restriction, and alternative sparse kernel-based models (e.g., Support Vector Regression (drucker1996support), One-Class SVMs (scholkopf2001estimating), Incremental SVMs (cauwenberghs2000incremental), or RVMs (tipping2001sparse)) can be used to extend DualXDA to regression, outlier detection, active learning, or provide probabilistic formulations.
Furthermore, the evaluation of DualXDA in the present study is limited to image classification models of moderate scale, and does not yet include a comprehensive quantitative assessment on contemporary large-scale or multimodal foundation models beyond the exploratory analyses provided in Appendix I. The method’s reliance on penultimate-layer representations suggests compatibility with other network families, including transformers. Our preliminary observations in Appendix I suggest that DualXDA may produce meaningful attributions for large transformer models in specific scenarios, although a systematic evaluation across model types, tasks, and data modalities remains to be conducted in future work.
The evaluation of XDA currently remains qualitative in nature. While the visualized attribution produced by XDA provide insight into feature-level interactions between training and test samples, quantitative evaluation metrics for such explanations are not yet established. Existing metrics for attribution map assessment have been developed for feature-attribution settings (hedstrom2023quantus) and do not directly apply to feature level attributions revealing interactions between training and test datapoints (instead of explaining classifier outputs), making the development of suitable evaluation measures an important direction for future work.
We note that the training of surrogate SVM models can introduce computational overhead for large training sets, although practical efficiency can be improved using early stopping or GPU-accelerated SVM implementations.
Finally, prior research suggests that combining Data Attribution methods over an ensemble of models may improve attribution efficacy. Considering the computational efficiency of DualXDA, this approach provides another potential avenue for extending DualXDA.
@@ 7 Discussion & Conclusion
In this work, we introduce DualXDA, a framework that enables efficient and sparse data attributions, coupled with feature-level explanations which clarify how specific training samples influence the prediction for a given test point. DualXDA consists of two parts: First, DualDA is an efficient surrogate-based DA method for explaining neural networks using the penultimate hidden representations to define a kernel map to fit a multiclass SVM. This strategy naturally provides global as well as local attributions as an approximate decomposition of the model output. Additionally, DualDA produces sparse attributions by virtue of the associated optimization process, which is beneficial for reducing the explanations’ complexity and volume. Second, XDA combines data attribution with feature attribution by mapping data attribution values back to the train and test input domain simultaneously, to identify relevant corresponding features and explain previously opaque data attribution results. Together, our framework overcomes the computational inefficiency of popular DA approaches and allows the user to gain more fine-grained insights and novel analytical capabilities.
@@
Finally, we employ XDA to gain deeper insight into the DualDA attributions across three different datasets. Through qualitative analysis, we demonstrate how XDA explains nuances in model predictions and confounding factors in training data, while showing robustness in input ablation studies. To our knowledge, this is the first work to propose a methodology to explain the interaction of training and test datapoints in the inference context of a trained model from both perspectives, on feature level. By unifying the attributions in two orthogonal approaches in XAI, feature attribution and data attribution, we obtain a method that leverages the benefits of both approaches, constituting a novel explanation paradigm.
@@ Appendix A Derivation of DualDA using the SVM framework
Given the soft-margin optimization problem of (svm:CS)
@@
To enforce the stationarity condition (svm:KKT), we take derivatives with respect to the primal variables and set them to 0:
@@
To solve this problem, we first derive the dual problem by plugging Equation A.8 in the Lagrangian and maximize it subject to the dual feasibility and complementary slackness conditions (svm:KKT):
We refer the reader to (svm:CS) for the derivations.
@@ Appendix C Derivation of DualDA using the Influence Functions Framework – Proof of Equation 8
For a given test point with features and an arbitrary class , our objective is to determine how the class score responds to the infinitesimal downweighting of the -th datapoint. Using Corollary 1 in (naujoks2024pinnfluence), we derive
@@ D.1 Data Attribution Methods
As DualDA is already defined in Section 3, we now specify the other DA methods we compare our approoach to in this paper. In all experiments, we employ GPU implementations and follow the best practices as suggested by original authors of the papers or code bases, as explained in the individual sections. This includes the use of random projections and CUDA-level implementations to efficiently perform these projections for TRAK. For the caching phase of DualDA, the extraction of latent features is run using GPU capabilities, while the training of the SVM uses CPU. The sequential nature of the underlying SVM optimization algorithm does not allow for a straightforward GPU parallelization of existing implementations (keerthi2008sequential; chiang2016parallel).
@@ D.1.1 Influence Functions
Influence Functions (IF) originates from the field of robust statistics, where it was initially used for the analysis of the effects of outliers on linear models (misc:influence_outlier; hammoudeh2022identifying). Koh and Liang (method:koh_influence) proposed using this method for local data attribution of deep neural networks. It approximates the effect of discarding a training point using a second-order Taylor approximation. However, this approximation is conditioned on the current trained parameters of the network and hence includes information about the decision making strategies employed by the trained model. Data attribution by IF is given by
@@
Computation of the inverse Hessian is computationally prohibitive for modern architectures with millions of parameters. For this reason, (method:koh_influence) propose to use the LiSSA algorithm (misc:influence_agarwal) which approximates the inverse Hessian vector product for each test point using an iterative procedure. While this mitigates the computational cost, it does not entirely solve the problem: computing explanations for a single test point can still take a long time, possibly in the order of hours (hammoudeh2022identifying). In our experiments, we have followed the authors’ suggestions to determine hyperparameters which produce meaningful influence values while minimizing the required computation time. First, we have used only the final layer of deep neural networks to compute gradients for influence estimation with LiSSA. Secondly, we have set the total number of iterations dedicated to the estimation of the inverse Hessian vector product to the training dataset size in each experiment.
@@
Finally, the authors of Grosse et al. (grosse2023studying) propose to solve the inverse-Hessian-vector product using an Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC), which is derived from an approximation of the Fisher information of the network. As hyperparameters, we have used the default values given in the original code release.
@@ D.1.2 TracIn
In the implementation of these methods, we have employed random projections on gradients to be able to cache training gradients, following (method:trackin). This allowed the attributions to be achieved in a practical time scale. We have used 128 random projections.
@@ D.2 Model Training
For CIFAR-10, we train a ResNet-18 (arch:resnet-he) with random cropping and flipping as data augmentations, with a learning rate of 0.0003. We further include a weight decay term with a coefficient of 0.01 as part of the loss term.
@@ E.1 Sanity Checks
Hanawa et al. (evaluation:hanawa) suggest two sanity checks to determine the quality of DA methods.
@@
Hanawa et al. further suggest that attributions should be able to detect different subpopulations in the same class (evaluation:hanawa). To create this condition, they propose using an arbitrary partitioning of classes, and group each set in the partition into a super group. In practice, we have selected pairs of classes to group. These pairs of classes are chosen to be dissimilar to ensure that different strategies are learned for samples belonging to different subclasses. After training a neural network on the modified dataset, a successful DA methods should attribute highly to training samples that belong to the same subclass as the original class of the test sample.
To express this condition formally, following (evaluation:hanawa), let be an oracle function that returns the true (sub)class of the datapoint given as its input. Similar to the Identical Class Test, we report
@@ E.2 Metrics Emulating Downstream Tasks
The following metric mimics the downstream task of identifying mislabeled samples in the training data (method:koh_influence; method:representer; method:trackin). Correctly labelled datapoints can be classified by a model by relying on similar training samples of the same class. In contrast, mislabeled samples diverge from the typical feature patterns of their labeled class and consequently require individual memorization rather than generalization. Therefore, mislabelled datapoints are highly influential on their own model prediction, allowing the user to leverage DA approaches to detect mislabeled datapoints. To simulate this scenario, we randomly change the labels of a subset of the training data and retrain the model. We then use the attribution method to calculate each training sample’s self-influence, . To quantify the performance of the mislabeling detection by a DA method, we traverse the training data set in descending order of self-influence and record the cumulative density function of the percentage of mislabeled samples we have identified. We report the area under the curve of this cumulative density function, linearly transformed to a score between 0 and 1, where 0 corresponds to the minimum achievable area, 1 represents the maximum achievable area, and random attribution by sampling from a uniform distribution yields an expected score of 0.5.
@@
Another candidate downstream application for DA is to identify the causes of shortcut learning and backdoor poisoning attacks. If we modify the training dataset to introduce an artificial shortcut to trigger a certain model prediction, DA methods should attribute the responsible training images highly whenever the shortcut feature is present in the test sample. Similarly to previous work (method:koh_influence; hammoudeh2022identifying), we formulate a shortcut detection metric: determine a perturbation which takes a sample and adds a shortcut perturbation on it. Importantly, applies a consistent and detectable perturbation to all chosen samples. We use a small box in the center, and a frame around the images in our experiments.
We select a class as the shortcut class, and randomly sample a subset of datapoints from that class to apply . By retraining the model on this perturbed training data, the model learns a connection between the artifact and the class . We now create a particular dataset for evaluation by applying the perturbation to all samples and collecting all perturbed samples, which are not originally from class but who, after the perturbation is applied, are classified as class , i.e. . This ensures that if a test point is now classified as class , it is primarily due to the influence of the perturbation and therefore, the perturbed training samples should be most influential for the classification as class . For all such test points, we report the Area Under the Precision-Recall Curve (AUPRC) score for the detection of perturbed training samples, following (hammoudeh2022identifying).
@@ E.3 Retraining-based Metrics
The counterfactual relevance of training data can be empirically assessed through model retraining under an identical training regime with modified training sets. As explained in Section 2.2, this approach is the motivation for methods based on Influence Functions (method:koh_influence). The subsequent metrics evaluate the correspondence between observed changes in model behaviour and the attributions of the training data.
@@
LDS measures the effectiveness of attributions as an estimator for the effect of leaving out training samples by retraining the model on general subsets of the training data (misc:TRAK). Given attributions , and a subset of training data indices , we define
@@
A downstream application of data attribution is dataset distillation: limiting the training dataset to fewer examples, while maintaining high test accuracy (guo2022deepcore; joaquin2024in2core). In contrast to LDS, we are not interested in the model performance for randomly chosen subsets of the training data, but instead for the highest attributed subset. To measure this in a metric, we need a strategy to determine the global importance of a training point for the model performance on the entire test set. To achieve this, we compute the average value of the absolute attributions of a training point over the test set. We then restrict the training data to the most globally relevant datapoints and measure the test loss of a model retrained on the restricted training set. We report the test loss when retrained on 10% of the data as well as a weighted average
@@
Inspired by previous work, which investigates the sensitivity of model predictions to ablation of the training data (ilyas2022datamodels), we measure the test loss of a model that was retrained on a train set which had its most influential datapoints removed. Similarly to Coreset Selection, we consider the highest attributed subset. But in contrast, we aim to judge how much exclusive relevant information this subset holds that would induce an increase in the loss when the subset is removed. We use the same strategy as in Coreset Selection to determine the global importance of training samples. We report the test loss when retrained without the 10% most influential data as well as a weighted average
@@ F.3 DualDA: Effect of sparsity hyperparameter
These analyses of the sparsity and accuracy of the surrogate SVM models, alongside corresponding evaluation results indicate that the choice of yields a sparse surrogate while preserving strong fidelity to the original model and stable attribution performance in the image classification task with convolutional neural networks. However, under alternative setups, where the task, data modality and model architecture differ, the optimal hyperparameter might change. In such cases, an independent hyperparameter optimization process may be necessary under these evaluation criteria. This process is considerable more tractable compared to competing approaches, due to the computational efficiency provided by DualDA.
@@ F.6 Details of Mislabeling Detection
F.7 Details of Sparsity Experiment
In Figure 5, we analyze the distribution of absolute attribution scores over the training set for various DA methods, exemplified on the AwA2 dataset. Since positive and negative attributions carry distinct semantic interpretations, it is worthwhile to investigate the distribution of positive and negative attributions for different method. In this appendix, we provide a more detailed view into the attributions, taking into account the distribution of negative and positive attributions. Initially, we examine the overall ratio of cumulative positive attribution versus negative attribution, displayed on the left side of the figure. Positive attributions show that a training point carries information corroborating the model prediction being explained, as such, they are expected to be concentrated on fewer datapoints from the corresponding class, compared to negative attributions which carry information about different kinds of evidence contradicting the prediction at hand. As expected, most explainers – with the exception of GradDot and GradCos – exhibit greater positive than negative cumulative attribution.
The proportion of negative to positive cumulative attribution varies significantly across methods. The LiSSA implementation of Influence Functions shows the lowest ratio at only 53.7%, while TRAK demonstrates the highest at 98.7%.
For DualDA, we observe that the sparsity hyperparameter directly affects this ratio – increasing values of (i.e., increasing sparsity) bring the positive cumulative relevance increasingly closer to the negative cumulative relevance.
On the right side of Figure F.9, we analyze the growth of the cumulative sum of both positive and negative attribution as a function of the fraction of most influential datapoints included, to assess the sparsity characteristics of attributions with distinct semantics. When comparing the development of the cumulative sum between positive and negative attribution, we observe that DualDA maintains consistent sparsity across both types. In contrast, Arnoldi and EK-FAC show increased sparsity for negative attribution, while Representer Points and GradDot demonstrate significantly reduced sparsity for negative attribution compared to positive attribution.